The condensed form of Nordstrop Company’s flexible budget for manufacturing overhead follows:

Cost Machine-Hours
Formula
Overhead Costs (per machine-hour) 8,500 9,500 10,500
  Variable cost $ 1.05 $ 8,925   $ 9,975   $ 11,025  
  Fixed cost 28,900   28,900   28,900  
  Total overhead cost $ 37,825   $ 38,875   $ 39,925  

The following information is available for a recent period:

  1. The denominator activity of 8,500 machine-hours was chosen to compute the predetermined overhead rate.
  2. At the 8,500 standard machine-hours level of activity, the company should produce 2,125 units of product.
  3. The company’s actual operating results were as follows:
  
  Number of units produced 2,425  
  Actual machine-hours 8,500  
  Actual variable overhead costs $ 9,800  
  Actual fixed overhead costs $ 29,000  

Required:

1. Compute the predetermined overhead rate and break it down into variable and fixed cost elements. (Round your answers to 2 decimal places.)

2. What were the standard hours allowed for the year’s actual output? (Do not round intermediate calculations.)

3. Compute the variable overhead spending and efficiency variances and the fixed overhead budget and volume variances. (Do not round intermediate calculations. Round your answers to 2 decimal places. Indicate the effect of each variance by selecting "F" for favourable, "U" for unfavourable, and "None" for no effect (i.e., zero variance).)
Answer #1
Part-a
Computation of Predetermin Overhead Rate
predetermine Overhead Rate= Estimated cost/ Machine Hour
37825/8500= $4.45
Variable Production Overhead Rate= Variable cost/ Machien Hour
8925/8500= $1.05
Fixed Production Overhead Rate= Fixed cost/ Machien Hour
=28900/8500= $3.40
Part-b
Standard Hour Allowed= Actual unit Produced X Std Machine Hour Require per Unit
2425 UnitX (8500/2125)= 9700 Hour
Part-c
Variable Overhead Efficiency Variance= ( Std Hou- Actual Hour) X Std Rate
=(9700-8500)*1.05= $1260 Favourable
Variable Overhead Rate Variance= ( Std Rate- Actual Rate)X Actual Machine Hour
=(1.05-(9800/8500))*8500= $875 Unfav
Fixed Overhead Budget Variance=Budgeted Fixed Overhead- Actual Overhead
28900-29000= $100 Unfavourable
Fixed Overhead Volume Variance= Std Fixed Overhead - Budgeted Fixed Overhead
=(8500*3.4)-28900= 0